Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
zaza - 16:01 Wednesday 19 November 2025 (68190) Print this report
Operator Report - Morning shift

7:00 UTC ITF found in DOWN with the following issues :

  • DRMI_LOCK node stuck due to SR misalignment process malfunction. Solved by acting on SUS_SR and SUS_PR nodes (operator)
  • Various error messages on different VPM processes (Masserot)
    • frame 1447574578 not send to dy_xxxxxxxx: sending queue is full
    • FdTOTProcessOne: 1447574000 nData error for V1:Daq_FbTrend100s_1000_dirRemovedGB
    • FbtBuildTrendFrame. Vector overflow at 1447438000 index=195 samplingRate=1 channel=V1:SIB2_B2_QD1_H_DC_missing serData:V1:SIB2_B2
    • etc ...

COMMISSIONING mode set (keeping LOCKED_ARMS_IR)

7:04 UTC Ncal rotor position metrology (Mours, INF)
8:15 UTC Measurements and door installation (partial) at MCB, WE, NE, CB (Fabozzi, Romboli)
9:37 UTC Neovan chiller assessment (Gherardini, Gosselin, De Rossi)


 

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×